# Fait: Fault-Aware Fine-Tuning for Better Code Generation

### v1

@@ 5. Discussion
+Fait | 82.9 (79.3) | 83.1 (70.6) | 48.9 | 20.3 |
Compared to Reinforcement Learning Method. Given the increasing popularity of reinforcement learning (RL) methods for improving code generation (Le et al., 2022; Wang et al., 2022), we believe it’s important to compare our approach with these established techniques. As RL methods in code generation typically aim to align model outputs with desired code solutions by increasing the probability of correct implementations while reducing the likelihood of erroneous ones, they share similarities with our work principle of enhancing LLMs’ ability to identify error-sensitive segments to improve code generation capabilities. Specifically, we compare our approach with the representative DPO method (Rafailov et al., 2023), which is widely used and has demonstrated significant advantages in code generation (Yang et al., 2024; Miao et al., 2024; Gee et al., 2024; Zhang et al., 2024; Gallego, 2024). This method works by training models to directly maximize the likelihood of preferred outputs over non-preferred ones without requiring explicit reward modeling, learning from paired examples of more and less desirable code implementations.
To ensure a fair comparison, we select the same LLMs and use identical experimental settings as stated in RQ1 for DPO training. The training dataset remains consistent across both DPO and Fait, and we evaluate and compare the performance of DPO and Fait on the Humaneval(+), MBPP(+), and BigCodeBench benchmarks.
Table 5 shows the performance of LLMs trained with different methods. Overall, LLMs with Fait consistently outperform those trained with DPO. We can observe that Fait outperforms DPO by a relative average of 4.2%, across three selected benchmarks. This advantage stems from fundamental methodological differences: while DPO relies on coarse-grained preference signals that cannot precisely target error-sensitive segments, Fait specifically maintains learning across all tokens while strategically emphasizing error-sensitive segments within code implementations. This approach ensures the LLM retains general coding knowledge while becoming more attentive to critical details that often determine functional correctness.
@@
6. Threats To Validity
There are three major threats to the validity of our work.
@@ 6. Threats To Validity
Threats to internal validity involve the impact of the quality of incorrect code and choices of hyperparameters. The effectiveness of our approach depends on the quality of incorrect code variants and the weighting factor . To mitigate threats related to code quality, we prompt a strong teacher model (Qwen2.5-Coder-32B-Instruct) to generate plausible incorrect variants and manually verify whether the generated data meets our expectations. While a small portion of noise data - incorrect implementations that differ completely from correct solutions - remains present, we argue these instances may actually enhance model robustness by preventing overfitting to specific error-sensitive segments (DeVries and Taylor, 2017; Xie et al., 2020). For hyperparameter-related threats, we conduct extensive sensitivity analysis as shown in Figure 7. In future work, we intend to investigate the use of stronger teacher models, such as GPT-4-Turbo, to generate similar incorrect code and examine their impact.
Threats to construct validity relate to the reliability of evaluation metrics. We evaluate our approach using the pass rates metric; however, this metric may inadequately capture the functional correctness of generated code implementations with a limited number of test cases. To address this limitation, we deliberately incorporate the extended versions of some benchmarks, which substantially expand the number of test cases. In future work, we plan to explore additional evaluation approaches, such as LLM-as-a-Judge (Gu et al., 2024), to provide a more comprehensive assessment of code quality beyond functional correctness.
7. Related Work
7.1. LLMs for Code Generation
@@
As a momentous milestone, Codex (Chen et al., 2021) boasting a 12-billion-parameter model demonstrates the extraordinary capability to tackle up to 72% of Python programming problems. After that, a new wave of code generation models, such as AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., [n. d.]), InCoder (Fried et al., [n. d.]) and StarCoder (Li et al., 2023) are proposed and have shown promising results in the code generation task. Building upon these foundations, more code-focused LLMs emerged, such as Magicoder (Wei et al., 2024), SemCoder (Ding et al., 2024), WaveCoder (Yu et al., 2024) and WizardCoder (Luo et al., [n. d.]). These specialized LLMs are typically based on general LLMs in solving domain-specific coding tasks through instruction tuning.
7.2. Fine-tuning on Code LLM
@@
Fine-tuning pre-trained language models has emerged as a dominant paradigm for optimizing performance in code generation. Instruction tuning (Ouyang et al., 2022; Achiam et al., 2023), as a form of supervised fine-tuning, aims to align LLMs with instruction through high-quality instruction corpora. For instance, Magicoder (Wei et al., 2024) introduce OSS-Instruct, an instruction tuning dataset generated by a teacher LLM drawing inspiration from open-source code snippets, which effectively enhances code generation capabilities when used for fine-tuning. Furthermore, OSS-Instruct is orthogonal to existing instruction tuning datasets like Evol-Instruct (Luo et al., [n. d.]), enabling the Magicoder series LLMs finetuned on this combined data to achieve further performance improvements. Similarly, SemCoder (Ding et al., 2024) propose PYX, a dataset created by a teacher LLM simulating human debugging processes. By incorporating data that simulates execution reasoning and captures code execution nuances, LLMs finetuned with PYX understand and articulate the execution process step-by-step, enhancing their reasoning capabilities. Likewise, combining with existing instruction tuning datasets like Evol-Instruct, the resulting SemCoder-S LLM further improves the original LLM’s code generation abilities.
To address limitations in preventing untruthful and unexpected outputs, researchers explore reinforcement learning (Ouyang et al., 2022). To address limitations in undesired outputs, researchers have explored reinforcement learning approaches. For instance, CodeRL (Le et al., 2022) utilizes compiler feedback as reward signals combined with REINFORCE (Williams, 1992) to fine-tune CodeT5 (Wang et al., 2021), reducing compilation errors. SENSE (Yang et al., 2024) synthesizes text-to-SQL training data from both strong and weak language models, and leverage Direct Preference Optimization (DPO) (Rafailov et al., 2023) to learn from correct and incorrect SQL examples, demonstrating state-of-the-art performance on text-to-SQL benchmarks. Similarly, PPOCoder (Shojaee et al., [n. d.]) trains CodeT5 with proximal policy optimization (Schulman et al., 2017). However, despite these advances, these fine-tuning approaches face a fundamental limitation: they treat all tokens with equal importance during loss calculation, making it difficult for models to distinguish semantically correct implementations from syntactically similar but incorrect ones. In this paper, we aim to address this challenge in code LLMs.
@@
We find recent work Focused-DPO (Zhang et al., 2025) enhances code generation capability by concentrating preference optimization on error-prone code points through an improved DPO (Rafailov et al., 2023) methodology. Our method is complementary to this approach - while Focused-DPO operates during post-training reinforcement learning stages, our approach operates during the supervised fine-tuning stage. Additionally, their work has demonstrated that their method can strengthen the capabilities of post-trained models. However, as Focused-DPO is currently under peer review and its implementation is not yet publicly available, we are unable to experimentally validate the potential synergies between our approaches in this work.
8. Conclusion
@@
In this paper, we introduce Fault-Aware Fine-Tuning (Fait), a novel fine-tuning technique that enhances code generation capabilities in instruction-tuned LLMs by refining their ability to distinguish between correct implementations and subtly incorrect variants. Unlike conventional supervised fine-tuning that treats all tokens equally, our approach identifies and prioritizes error-sensitive segments through two key components: error-sensitive segments identification, that captures both line-level and token-level critical differences, and dynamic importance reweighting that dynamically adjusts token weights during training to focus on discriminative elements. Through extensive experiments across seven LLMs and three widely-used benchmarks, we demonstrate that our method achieves an average relative improvement of 6.9% on pass@1 with just one epoch of training, with certain enhanced 6.7B LLMs even outperforming GPT-3.5-Turbo on selected benchmarks. The technique also exhibits strong generalization capabilities across diverse instruction-tuned LLMs and maintains effectiveness even when applied to LLMs with closed-source instruction datasets, providing a practical solution to improve LLMs’ code generation capabilities.

### v2

@@ V Discussion
+FGit | 82.9 (79.3) | 83.1 (70.6) | 48.9 | 20.3 |
Compared to Reinforcement Learning Method. Given the increasing popularity of reinforcement learning (RL) methods for improving code generation [42, 43], we believe it’s important to compare our approach with these established techniques. As RL methods in code generation typically aim to align model outputs with desired code solutions by increasing the probability of correct implementations while reducing the likelihood of erroneous ones, they share similarities with our work principle of enhancing LLMs’ ability to identify error-sensitive segments to improve code generation capabilities. Specifically, we compare our approach with the representative DPO method [44], which is widely used and has demonstrated significant advantages in code generation [45, 46, 47, 48, 49]. This method works by training models to directly maximize the likelihood of preferred outputs over non-preferred ones without requiring explicit reward modeling, learning from paired examples of more and less desirable code implementations.
To ensure a fair comparison, we select the same LLMs and use identical experimental settings as stated in RQ1 for DPO training. The training dataset remains consistent across both DPO and FGit, and we evaluate and compare the performance of DPO and FGit on HumanEval(+), MBPP(+), and BigCodeBench. Table V shows the performance of LLMs trained with different methods. Overall, LLMs with FGit consistently outperform those trained with DPO. We can observe that FGit outperforms DPO by a relative average of 4.2%, across three selected benchmarks. This advantage stems from fundamental methodological differences: while DPO relies on coarse-grained preference signals that cannot precisely target error-sensitive segments, FGit specifically maintains learning across all tokens while strategically emphasizing error-sensitive segments within code implementations. This approach ensures the LLM retains general coding knowledge while becoming more attentive to critical details that often determine functional correctness.
@@
Compared to Mutation-Based Method
Method | Humaneval (+) | MBPP(+) |
Avg | ||||
Original | 79.3 (74.4) | 79.6 (68.5) | 48.5 | 16.9 | 61.2 | ||||
Mutation-Based | 81.1 (77.4) | 80.9 (67.7) | 49.2 | 17.5 | 62.3 | ||||
FGIT | 83.5 (78.7) | 83.1 (70.6) | 48.9 | 20.3 | 64.2 |
Traditional mutation-based approaches provide a natural baseline for our method, as they are also designed to introduce errors into code. We use universalmutator [50], a state-of-the-art, language-agnostic, rule-based mutation tool, to generate faulty versions of the code in Evol-Instruct to compare with FGit. We train SemCoder-S using these mutated samples, keeping all hyperparameters consistent. As shown in Table VI, FGit outperforms the mutation-based approach. We attribute this to the fact that LLMs, trained on vast code corpora, generate a more diverse and realistic spectrum of plausible errors than a fixed set of mutation rules, leading to more effective learning.
Potential Overfitting of Standard-SFT
Model | Dataset | Humaneval (+) | MBPP (+) |
Avg | ||||
MagiCoder-DS | None | 66.5 (60.4) | 75.4 (61.9) | 43.4 | 12.2 | 53.3 | ||||
OSS-Instruct (Original) | 64.6 (58.5) | 79.1 (66.1) | 43.9 | 12.2 | 54.1 | |||||
PYX (New) | 66.5 (60.9) | 77.5 (65.8) | 45.7 | 15.5 | 55.3 | |||||
Evol-Instruct (New) | 64.6 (60.9) | 77.2 (65.6) | 47.3 | 15.5 | 55.2 | |||||
SemCoder | None | 73.2 (68.9) | 79.9 (65.3) | 43.5 | 16.9 | 58 | ||||
PYX (Original) | 71.3 (65.2) | 79.9 (66.5) | 43.4 | 14.2 | 56.8 | |||||
OSS-Instruct (New) | 73.2 (69.5) | 78.0 (64.3) | 46.4 | 18.9 | 58.4 | |||||
Evol-Instruct (New) | 71.3 (67.6) | 78.6 (65.6) | 47.6 | 19.6 | 58.4 |
We observe that Standard-SFT occasionally lead to performance degradation compared to the base model. For instance, as detailed in Table I, MagiCoderS-DS exhibit a decline in performance across multiple benchmarks after undergoing Standard-SFT. We hypothesize that this can be attributable to potential overfitting, where the model is trained on a dataset whose data distribution is already familiar from its initial fine-tuning phase. To validate this, we fine-tune models on datasets they have not previously been exposed to. Specifically, we fine-tune MagiCoder-DS on PYX and Evol-Instruct, and SemCoder on OSS-Instruct and Evol-Instruct. As shown in Table VII, the models achieve more gains when trained on unseen datasets. For example, SemCoder, when fine-tuned on either OSS-Instruct or Evol-Instruct, outperforms its results when trained on its original dataset, PYX.
VI Threats To Validity
@@
Threats to internal validity involve the impact of the quality of incorrect code and choices of hyperparameters. The effectiveness of our approach depends on the quality of incorrect code variants, the weighting factor and training epochs. To mitigate threats related to code quality, we prompt a strong teacher model to generate plausible incorrect variants. Subsequently, we manually and quantitatively analyze the portion of generated data which is incorrect yet similar to the correct solutions. While a small portion of noise data remains present, we argue these instances may actually enhance model robustness by preventing overfitting to specific error-sensitive segments [51, 52]. For hyperparameter-related threats, we conduct an extensive sensitivity analysis as shown in Figure 7. In future work, we intend to investigate the use of stronger teacher models, such as GPT-4-Turbo, to generate similar incorrect code and examine their impact.
VII Related Work
VII-A LLMs for Code Generation
@@
As a momentous milestone, Codex [6] boasting a 12-billion-parameter model demonstrates the extraordinary capability to tackle up to 72% of Python programming problems. After that, a new wave of code generation models, such as AlphaCode [53], CodeGen [54], InCoder [55] and StarCoder [56] are proposed and have shown promising results in the code generation task. Building upon these foundations, more code-focused LLMs emerged, such as Magicoder [2], SemCoder [18], WaveCoder [5] and WizardCoder [1]. These specialized LLMs are typically based on general LLMs in solving domain-specific coding tasks through instruction tuning.
VII-B Fine-tuning on Code LLM
@@
Fine-tuning pre-trained language models has emerged as a dominant paradigm for optimizing performance in code generation. Instruction tuning [57, 58], as a form of supervised fine-tuning, aims to align LLMs with instruction through high-quality instruction corpora. For instance, Magicoder [2] introduce OSS-Instruct, a dataset generated by a teacher LLM drawing inspiration from open-source code snippets, which effectively enhances code generation capabilities. Similarly, SemCoder [18] propose PYX, a dataset created by a teacher LLM simulating human debugging processes. By incorporating data that simulates execution reasoning and captures code execution nuances, LLMs finetuned with PYX understand and articulate the execution process step-by-step, enhancing their reasoning capabilities. To address limitations in preventing untruthful and unexpected outputs, researchers explore reinforcement learning [57]. To address limitations in undesired outputs, researchers have explored reinforcement learning approaches using preference optimization [45, 59], like DPO [44], to refine outputs. However, these often treat tokens uniformly in their loss calculations, making it difficult for models to distinguish semantically correct implementations from syntactically similar but incorrect ones. In this paper, we aim to address this challenge in code LLMs.
@@
We find recent work Focused-DPO [60] enhances code generation capability by concentrating preference optimization on error-prone points through an improved DPO [44] methodology. Our method is complementary to this approach - while Focused-DPO operates during post-training reinforcement learning stages, our approach operates during the supervised fine-tuning stage. However, as Focused-DPO is under peer review and its implementation is not yet publicly available, we are unable to experimentally validate the potential synergies between our approaches in this work.
VIII Conclusion
@@
In this paper, we introduce Fault-Guided Fine-Tuning (FGit), a novel fine-tuning technique that enhances code generation capabilities in instruction-tuned LLMs by refining their ability to distinguish between correct implementations and subtly incorrect variants. Through extensive experiments across seven LLMs and three widely-used benchmarks, we demonstrate that our method achieves an average relative improvement of 6.9% on pass@1, with certain enhanced 6.7B LLMs even outperforming GPT-3.5-Turbo on selected benchmarks. The technique also exhibits strong generalization capabilities across diverse instruction-tuned LLMs and maintains effectiveness even when applied to LLMs with closed-source instruction datasets.